AO Com Work: 25 October 2010 UT

Observer: DMiller, SEsposito, MXompero, LBusoni, APuglisi, FQuiros
Telescope Operator: DHuerta (LBTO)
Telescope Support: MMidkiff (LBTO)
Instrument Support: TShih (LBTO)

Summary

Opened with high cirrus clouds, 50% humidity, 12 m/s wind

Preparations

Details

01:00 Opened with high cirrus clouds. David set IE=-22 CA=37

01:30 Preset ACTIVE to AO_20p40m09_1. No star found. Removed this catalog from svn

01:35 Preset ACTIVE to AO_20p40m09_0. Dim because of clouds. AO guys close the loop and do some work

02:10 Preset ACTIVE to AO_22p20m08_0 to the South East. Collimate then loose star to clouds

02:30 Preset ACTIVE to AO_22p20m08_0 and acquired star. A bit brighter

Seeing is really bad: 1.5" to 4" on the DIMM.

AO guys are able to close the loop, but not collect much data worth much

04:00 point at IO. Close the loop in 1.5 to 3.0" seeing. Offset, using the engineering GUI, so Jupiter visible in the IRTC. First time on-sky to collect data using the W-Unit off-axis.

05:30 Close due to rising humidity

06:00 Open for 15 minutes and then high humidity

07:13 Open. preset ACQUIRE M5_1040

07:25 Set IE=-41, CA=40. Collimated on AO_02p40m08_2, gs=2. Seeing is 2-3".

07:35 AO guys close the loop

07:45 Closed for high humidity

-- DougMiller - 25 Oct 2010
Topic revision: r3 - 27 Oct 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback